OpenStudy (anonymous):

A golfer hits a ball from ground level with an initial vertical velocity of 64 ft/s. After how many seconds will the ball hit the ground?

OpenStudy (anonymous):

starting at ground level initial position = y0 = 0, final position = y = 0, gravity = a = -9.80, gravity is negative because it point down in the x-y coordinate system

OpenStudy (anonymous):

y= y0 + v0*t + 1/2 a*t^2 v0 = initial velocity

OpenStudy (anonymous):

y-y0 =v0*t +1/2 *a *t^2 rearrange so it is easier to see it is a quadratic equation 1/2* a*t^2 +vo*t + (y-y0) 1/2*a =- 4.9 v0= 64 y-y0 = 0 -4.9*t^2 +64*t +0 = 0 plug into quadratic equation and you get t = 13.06122448 t = 0 t = 0 don't make since so we use 13.06

OpenStudy (anonymous):

sorry had to make a correction * -4.9

OpenStudy (anonymous):

another correction * 64 ft /s have to be covert to m/s to have same units a gravity

OpenStudy (anonymous):

apply 2nd eq of motion here only